The area of a rectangle is 1,872 ft2. The ratio of the length to the width is 9:13. Find the perimeter of the rectangle.

Answer by oscargut(2103)

Let x = width and y = length
x/y = 9/13
xy = 1872
x = 9y/13
9y^2/13 = 1872
y^2 = 2704
y = 52
x = 36
Perimeter = 2(36+52) = 176
If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks